A proton and an electron are initially at rest at a distance of 9×10-¹⁰m. What will be their initial acceleration due to the electric force that they exert on each other?

1
Expert's answer
2022-02-21T17:04:02-0500

"F=\frac{kq_1q_2}{r^2}" "=\frac{9×10^9(1.6×10^{-19})^2}{(9×10^{-10})^2}"

"=2.844\u00d710^{-10}"


"F=ma"


Acceleration of electron

"a=\frac{2.844×10^{-10}}{9.11×10^{-31}}"


"=3.117\u00d710^{20}m\/s^2"



Acceleration of proton


"a=\frac{2.844×10^{-10}}{1.67×10^{-27}}"


"=1.703\u00d710^{17}m\/s^2"